OpenStudy (anonymous):

Rewrite with only sin (x) and cos (x): cos (2x) - sin (x) Can anyone help with this? Please I am so lost and it's been a while since I've worked on pre-calc

OpenStudy (anonymous):

there are a couple ways to do this

OpenStudy (anonymous):

they all involve using the double angle formula for cosine, but there are three ways to write it \[\cos(2x)=\cos^2(x)-\sin(2(x)\] \[\cos(2x)=2\cos^2(x)-1\] \[\cos(2x)=1-2\sin^2(x)\] any one of those will work